OpenStudy (anonymous):

to what temperature will a 50.0 g piee of glass raise if it absorbs 5275 joules of heat and its specific heat capacity is 0.50 J/g ºC? The initial temperature of the glass is 20.0ºC

OpenStudy (anonymous):

This is a thermal physics problem involving the popular equation: \[Q=mc \Delta T\] Q = 5275 J m = 50.0 g = 0.0500 kg c = 0.50 J/g degrees C Ti = 20.0 degrees C Tf = ? degrees C \[Q=mc \Delta T=mc(Tf-Ti)\]\[\frac{ Q }{ mc }+Ti=Tf=\frac{ (5275) }{ (0.0500)(500.0) }+(20.0)=231\] The glass will rise to a temperature of 231 degrees Celsius.